OpenStudy (anonymous):

Company "A" sells soil at 11cents per cubic feet. While company "B" sells it at 3 dollars per cubic yard. You need to cover 20 cubic yards. What is the best company? and how much will you save?

OpenStudy (shane_b):

All you need to know is that \[1 yd^3 = 27ft^3\]...from there you should be able to see which would be cheaper.

OpenStudy (anonymous):

I think B

OpenStudy (anonymous):

im just guessing though

OpenStudy (shane_b):

The answer is whichever is smaller:\[\frac{300c}{27ft^3}\]or\[\frac{11c}{1ft^3}\]

OpenStudy (anonymous):

So 11c/1ft3 Is cheaper its 11 while the other one is 11.11111etc. So is it A, or B. and how much do you save?

OpenStudy (anonymous):

So is that company A or B?

OpenStudy (anonymous):

If B i deserve a medal for being straight up

OpenStudy (shane_b):

lol

OpenStudy (anonymous):

any day now

OpenStudy (shane_b):

Company A is cheaper by 0.11c per cubic foot. You need 20 cubic yards so converting that to cubic feet you get: \[(20)(27) = 540 ft^3\] And you save:\[(540ft^3)(0.11)=$59.40\]

OpenStudy (shane_b):

crap.

OpenStudy (anonymous):

But on my math it says do you save, 40 or 60 cents?

OpenStudy (shane_b):

That last part should be:\[(540ft^3)(0.0011c)=59.40c\] They took it out to more places (0.0011111111...) than I did and rounded it to 60c.

OpenStudy (anonymous):

Ok, thanks:)

OpenStudy (shane_b):

Good luck :)

OpenStudy (anonymous):

It better be B

OpenStudy (shane_b):

BTW: @swagdaddy91: It was not Company B!

OpenStudy (anonymous):

.............

OpenStudy (anonymous):

@swagdaddy91 no worries your still the smartest person EVER!!!!!!!!!!

OpenStudy (anonymous):

You were right, If I need anything else. I will write it on here. I'm taking an online math course, cuz I don't want my Geometry teacher for next year. So I took this instead. It's really freaking hard -_-

OpenStudy (anonymous):

Shane B, You are buying tile to cover a floor. Half are blue, and Half are brown. Brown = 58 cents per sq.ft. and Blue = 49 cents per. sq.ft. How much would that be?

OpenStudy (shane_b):

Do you have more info than that?

OpenStudy (anonymous):

No. well the answers are. 30, 33.32, 39.44, and 72.26. What one is it?

OpenStudy (anonymous):

how big is the floor

OpenStudy (anonymous):

Oh sorry the floor is 17 by 8 sorry -_-

OpenStudy (shane_b):

that helps :)

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Swagdaddy91, can you leave because I NEED to get this done and you are very obnoxiousness .

OpenStudy (anonymous):

Is it 30?

OpenStudy (shane_b):

The total area will be 17*8=136ft^2 Half of that area is 68 So you have to buy 68 blue tiles @ 0.49c each and 68 brown tiles at 0.58c each. 68(0.49)+68(0.58)=73.25

OpenStudy (shane_b):

Oh..calculator mistake. 72.76

OpenStudy (shane_b):

68(0.49)+68(0.58)=72.26

OpenStudy (anonymous):

how am i obnoxious not my fault tht your taking a class online so you dont have to go to a certain teacher and then asking for answers for every single question you have.

OpenStudy (anonymous):

Well I kinda wanna pass, plus you had the WRONG answer so whatever.

OpenStudy (anonymous):

thts not being obnoxious

OpenStudy (anonymous):

It kinda wanna get it done, so yeah it is.

OpenStudy (anonymous):

nope you can get it done if you stop worrying about what im writing

OpenStudy (anonymous):

Shane, What is the area if a similar triangle; if triangle A has an height of 2 and a base of 5, and the other triangle has a base of 3. What is the area of B?

OpenStudy (anonymous):

My choices are 1, 1.8, 3.2, and 5?

OpenStudy (anonymous):

still cnt do your own work what a shame

OpenStudy (anonymous):

Then why are you on here? It's to get help for your example I have done over 13 hours of work on plato so STFU! I want help ok.

OpenStudy (anonymous):

lol u mad bro?